### v1

# Optimal Execution with Reinforcement Learning

Abstract
This study investigates the development of an optimal execution strategy through reinforcement learning, aiming to determine the most effective approach for traders to buy and sell inventory within a limited time frame. Our proposed model leverages input features derived from the current state of the limit order book.
To simulate this environment and overcome the limitations associated with relying on historical data, we utilize the multi-agent market simulator ABIDES, which provides a diverse range of depth levels within the limit order book.
We present a custom MDP formulation followed by the results of our methodology and benchmark the performance against standard execution strategies. Our findings suggest that the reinforcement learning-based approach demonstrates significant potential.
Keywords: Optimal Execution, Limit Order Book, Reinforcement Learning, Agent-Based Simulation, Algorithmic Trading, Transaction Costs, Market Impact.
I Introduction
In the intricate landscape of financial markets, optimizing the execution of significant positions remains a critical challenge for financial institutions such as banks, asset management firms, hedge funds, and prop shops.
Research in market microstructure shows that large trades influence asset prices because the immediate depth of the market is limited [Bouchaud et al., 2009]; a single large order can exhaust all current buyers or sellers. This suggests that it is generally advantageous to split large orders into several smaller blocks. Furthermore, nuances in portfolio adjustment can cause adverse price changes, forcing traders to strike a delicate balance between quick trading and possible poor execution, or slower trading and exposure to unpredictable market fluctuations [Gueant, 2016].
Various techniques are used to minimize market impact, which generally consist of splitting the large order into smaller ones. Traditional approaches, typified by the Almgren-Chriss approach [Almgren and Chriss, 2001], rely on stochastic optimal control to maximize performance and are solved analytically. For example, [Almgren, 2012] explored the crafting of optimal execution strategies in the context of varying market liquidity and volatility over time. [Cartea et al., 2018] developed a high-frequency trading strategy, using superior speed for information processing and order placement while introducing a multifactor mutually exciting process to enable feedback effects on market orders and the shape of the limit order book. Meanwhile, [Giacinto et al., 2022] examined optimal liquidation within a market containing various diverse market makers with constrained inventory-holding and risk-bearing capacities.
However, the approach to solving the stochastic control problem is limited. Analytical solutions are rare and require stringent conditions on the problem’s modeling. Other approaches involve solving or approximating the Hamilton-Jacobi-Bellman (HJB) equation (see [Bertsekas, 2005]) or Quasi-Variational inequalities (see [Carmona and Delarue, 2002, Øksendal and Sulem, 2007]), or even addressing the problem through its probabilistic formulation using backward stochastic differential equations (BSDEs) [El Karoui and Quenez, 1997]. These methods are constrained and often rely on viscosity solutions to prove the existence and uniqueness and characterize the regularity of the value function [Crandall and Lions, 1992]. Furthermore, numerical schemes to approximate these equations, such as Monte-Carlo methods (see [Glasserman, 2004]) for BSDEs and other discretization methods for PDEs, struggle with high-dimensional problems (beyond three to four dimensions), significantly limiting the modeling capability.
Reinforcement learning techniques have emerged as viable solutions, exploring sequential decision problems without stringent modeling assumptions. These methodologies, applied in solving the optimal execution problem provide avenues when analytical solutions are elusive.
One of the challenges in studying this problem is that, in historical simulations, it is not possible to reproduce the effect of the trade without making assumptions about market impact. Therefore, we address this issue using the multiagent market simulator ABIDES [Byrd et al., 2020] to train and test our approach. This approach mitigates the challenges encountered when studying this problem, especially in historical simulations, where accurately reproducing trade effects without assuming market impact is difficult. To stabilize the profit and loss (P&L) dynamics, we impose penalties on trade-throughs—instances where at least one limit in the order book is depleted. Additionally, we implement a delayed penalty reward to ensure that agents do not retain any remaining inventory at the end of the execution period.
II The Limit Order Book
Figure 1: Representation of LOB
A Limit Order Book (LOB) is a comprehensive record of current limit orders, where price changes occur discretely in increments known as the tick size (see Figure 1). Each order consists of a price and size, collectively contributing to the volume represented by $Q$. One of the key elements highlighted in the Limit Order Book is the concept of liquidity. In a perfectly liquid market, any quantity of a specific security can be immediately converted to cash. Measurement of liquidity involves metrics such as traded value / turnover, bid-ask spread, and trade-through intensity, as described in [Chevalier et al., 2023].
Order Types
When trading in a LOB, it is possible to employ different types of orders. The limit order is the most used and entails specifying the price at which we want to execute the trade. There is no guarantee that the trade will happen, as the limit price may never be reached. These orders are executed considering a price-time priority, also known as First-In-First-Out (FIFO). Then there is the market order, which consists of a request to carry out the order immediately at the best price available in the market. To be more specific, a buy (sell) order is matched with limit sell (buy) orders starting with the best ask price. Finally, there are other types of orders, such as fill-or-kill orders and others.
LOB features
In algorithmic trading, it is popular to use signals derived from specific characteristics, including the following:
  • Total depth: it corresponds to the cumulative sum of volume at multiple price levels ($d$) of the LOB:
[…]
    where $Q_{\text{ask}}^{k}$ ($Q_{\text{bid}}^{k}$) is the volume of outstanding limit orders at the $k$-th best ask (bid) price level.
  • Volume imbalance: it describes the difference between the existing order volume on the bid and ask price levels. It can be defined as:
[…]
  • Mid price: it is the midpoint between the best bid and best ask prices:
    $P_{\text{mid}}=\frac{P_{\text{best ask}}+P_{\text{best bid}}}{2}.$
  • Spread: it is the difference between the best ask and the best bid:
    $P_{\text{mid}}=P_{\text{best ask}}-P_{\text{best bid}}.$
II-A Modelling the LOB
Traditional LOB models often overlook key market features, such as heavy tails and volatility clustering. The early works [Stigler, 1964] and [Garman, 1976] focused on market regulations and microstructure using simple simulations. Later models introduced different types of trades and mechanisms, such as order flow and herding behavior (see [Gould et al., 2013]). These new approaches leverage large datasets and focus on better replication of various stylized facts such as heavy tails and volatility clustering.
Stochastic Models
Stochastic models (also referred to as Zero Intelligence models) represent the dynamics of the LOB using probabilistic processes. These models often describe the arrival of orders and their execution using Poisson processes or other random mechanisms. Stochastic models are valuable for their analytical tractability and the ability to derive closed-form expressions for various market metrics. The seminal work [Cont et al., 2010] provides a comprehensive stochastic model for the LOB, capturing key features such as order flow and market depth, and a fast estimation method from market data. Other notable examples include [Smith et al., 2003, Huang et al., 2013, Tommaso Mariotti and Toscano, 2023].
Machine Learning Approaches
Machine learning techniques, particularly deep learning, have gained popularity in LOB modeling due to their ability to capture complex, nonlinear patterns in data. These methods do not require agent calibration, but instead learn simulated market behavior directly from historical data. Generative Adversarial Networks (GANS) [Cont et al., 2023, Li et al., 2020, Coletta et al., 2022a] are commonly used to model temporal dependencies in LOB data.
Agent-Based Models
Agent-Based Models (ABMs) simulate the interactions of autonomous agents, each following a set of rules or strategies. These models are particularly useful for capturing the heterogeneous and strategic behavior of market participants. ABMs offer insight into individual agent performance and the overall impact of their interactions. These models lie between zero-intelligence and perfect-rationality models, as they permit agent behaviors to be specified without the need for explicit rationality. Examples of ABMs for LOB modeling include [Alfi et al., 2009, Chakraborti et al., 2011, Hamill and Gilbert, 2015, Coletta et al., 2022b].
Why ABIDES?
After reviewing various approaches, we decided to use the ABIDES (Agent-Based Interactive Discrete Event Simulation) [Byrd et al., 2020] framework for our research. ABIDES is designed to simulate realistic financial markets with high fidelity. It supports the creation of diverse agent behaviors, including those of institutional and retail traders, market makers, and high-frequency traders. ABIDES allows for detailed modeling of LOB dynamics and the interactions between different market participants.
II-B ABIDES Configuration
Let $(\Omega,\mathbb{F},\mathcal{F}=\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a complete filtered probability space endowed with right-continuous filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$. Going forward, whenever we mention equalities between random variables, we mean that they hold true almost surely under the probability measure $\mathbb{P}$. We may not explicitly indicate this notion ($\mathbb{P}$a.s.) in most cases. In the simulation, various agents, such as Exchange Agents, Adaptive Agents, Market Maker Agents, Value Agents, and Momentum Agents, develop their trading strategies based on an estimate of the fundamental value, which is defined by an Ornstein-Uhlenbeck (OU) process. The dynamics of the fundamental process process can be described by the following stochastic differential equation (SDE):
$dX_{t}=\theta(\mu-X_{t})\,dt+\sigma\,dW_{t}+J\,dN_{t},$ (1)
where:
  • $\theta$ is the rate of mean reversion,
  • $\mu$ is the long-term mean,
  • $\sigma$ is the volatility,
  • $W_{t}$ is a standard Wiener process,
  • $N_{t}$ is a Poisson process with intensity $\lambda$, representing the arrival of major news events,
  • $J$ is the jump size, drawn from a bimodal distribution centered at zero. Specifically, $J$ can take values from two Gaussian distributions: $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$, where $\mu_{1}=-\mu_{2}$ and $\sigma_{1}=\sigma_{2}$, ensuring the overall mean of the bimodal distribution is zero.
The jump process $N$ is introduced to represent events related to extrinsic news of substantial nature that occur relatively infrequently but have the potential to significantly alter the consensus valuation of a stock. This allows us to obtain improvements of the simulation computation time while producing more realistic price time series.
We chose the RMSC-4 configuration, which is the reference configuration in ABIDES and was also used in [Amrouni et al., 2022] and [Shi and Cartlidge, 2023]. Several key settings are defined within the RMSC-4 configuration to realistically simulate market behaviors and participant interactions. Exchange agents manage order books and historical data streams, maintaining a depth of 10 levels and a history of 500 streams.
Noise agents introduce randomness into the market through the actions of 1000 agents.
Value agents, numbering 102, base their trades on a true mean fundamental value, which behaves like an Ornstein-Uhlenbeck (OU) process centered around $\mu_{va}$ of \$100,000. This process has a mean reversion parameter $\theta_{va}$ set at $1.67\times 10^{-15}$ and an arrival rate $\lambda_{va}$ of $5.7\times 10^{-12}$.
An oracle-like agent tracks the mean-reversion process with a parameter $\theta_{or}$ set at $1.67\times 10^{-16}$ and a volatility of $5\times 10^{-10}$ for the fundamental time series, providing a benchmark for the fundamental value.
Market Maker agents dynamically adjust their pricing strategies. They use an adaptive window size, set their order size to $0.025$ percent of the volume, maintain a price range within 10 ticks, and operate with a wake-up frequency of $1$ second to update their quotes. This setup uses a higher frequency compared to other standard configurations, making it more realistic and better adapted to today’s markets (e.g. [Nagy et al., 2023, Karpe et al., 2020]).
Momentum agents, totaling $12$, trade according to recent market trends, reacting to short-term price movements.
These settings are designed to provide a comprehensive simulation of market dynamics, reflecting the diverse strategies and behaviors of market participants within a controlled environment.
Figure 2: Ask prices generated by ABIDES for different seeds.
III Optimal Execution Setting
The optimal execution problem consists in executing a trade of $X_{0}$ shares in a maximum amount of time $T$. In assessing the trade, it’s essential to consider its immediate transactional impact, temporary price fluctuations, and potential long-term effects on future prices due to lasting alterations in pricing dynamics. Additionally, a penalty could be considered if the agent fails to accomplish complete execution within the specified time period $[0,T]$. In what follows, we will consider an optimal liquidation problem without loss of generality.
In a discrete setting with a number of time-steps $N+1$, the execution strategy is a sequential decision process in which the trader decides the quantity to execute $x_{k}$ in each time step $t_{k}=k\frac{T}{N}$ for $k\in\{0,...,N\}$, where $t_{0}=0$ and $t_{N}=T$. A trading trajectory is defined as a list $\{x_{0},...,x_{N}\}$, where $x_{k}$ is the number of units held at time $t_{k}$. Clearly, $x_{0}=X_{0}$ and liquidation at time $T$ requires $x_{N}=0$, that is, $\sum_{k=0}^{N}x_{k}=X_{0}$. Let $P_{n}$ denote the average liquidation price for trade $x_{n}$ [^1] The execution price $P_{n}$ typically depends on both the current trade $x_{n}$ and all preceding trades. Additionally, we consider that $P_{n}$ encompasses all transaction costs.. The goal is to minimize the expected total liquidation cost [^2] This objective prioritizes expected value for a risk-neutral trader. It can be extended to include more risk factors. over the horizon $T$ :
$\min_{x\in\mathcal{A}}\mathbb{E}\left[\sum_{k=0}^{N}P_{k}x_{k}\right],$ (2)
$\mathcal{A}=\left\{\left\{x_{0},x_{1},\ldots,x_{N}\right\}\in\mathbb{R}^{N+1}_{+};\sum_{k=0}^{N}x_{k}=X_{0}\right\}.$ (3)
The most commonly used execution algorithm is the Time Weighted Average Price (TWAP). TWAP aims to execute a trade of size $X_{0}$ evenly over a specified time horizon $T$ in sizes of $X_{0}/N$. The arithmetic average of prices collected yields the TWAP price:
$\text{TWAP}=\frac{X_{0}}{N}\sum_{k=0}^{N}P_{k},$ (4)
where $P_{k}$ is the average at which each trade was executed through a market order sent at time $t_{k}$. In Almgren-Chriss terms [Almgren and Chriss, 2001], TWAP corresponds to a scenario with zero risk aversion, where the trader is indifferent to price volatility and only aims to minimize market impact by trading evenly.
IV Related Works
Initial studies on optimal execution are based on a stochastic framework in which the dynamics of the assets and execution costs are modeled through SDEs. The first formal analysis of the optimal execution problem is by [Bertsimas and Lo, 1998]. Under suitable conditions and using dynamic programming, they provide a closed-form formula as a solution to the optimal execution problem. Subsequently, an extension of the stochastic model was provided by introducing permanent and temporary effects due to the impacts of orders on the market and by inserting a risk-aversion parameter by [Almgren and Chriss, 2001]. The Almgren-Chriss model optimizes the trading strategy $x(t)$ over a time horizon $[0,T]$ to minimize total costs, and has a parameter to adjust the trade-off between minimizing market impact and execution costs.
[Huberman and Stanzl, 2005] extended the optimal execution framework by introducing more complex price impact functions and risk aversion parameters. Furthermore, the book by [Gueant, 2016] proposes an in-depth analysis and extension of these approaches. These methods, however, make strong assumptions on the underlying price movement or distributions.
Thanks to technological advances and data availability, [Nevmyvaka et al., 2006] applied RL for the first time to optimal execution strategies, with the objective of minimizing the implementation shortfall. Subsequently, [Hendricks and Wilcox, 2014] proposed to combine the Almgren-Chriss model with the Q-learning algorithm, to create a hybrid framework that executes a proportion of the AC trajectory based on the states in input. To address the high dimensions and complexity of the underlying dynamics, [Ning et al., 2018] used the Deep Q-Network (DQN) [Mnih et al., 2015], a combination of deep neural network and Q-learning, for optimal trade execution, addressing the curse of dimensionality issue faced by tabular Q-learning. [Lin and Beling, 2020], one of the most recent works, used PPO [Schulman et al., 2017a], to propose an optimal execution framework that can account for temporal correlations and make decisions based on LOB data using a sparse reward signal. These approaches suffer from the shortcomings of learning from historical data: the impossibility of reproducing realistically the impact of the orders of the agent.
This development has facilitated multiagent approaches [Balch et al., 2019] and led to the creation of ABIDES [Byrd et al., 2020]. [Karpe et al., 2020] was the first to utilize ABIDES to simulate a realistic trading environment and applied the DDQL algorithm (e.g. [Van Hasselt et al., 2016]) to learn optimal execution policies. Furthermore, [Nagy et al., 2023] formulated an execution strategy based on predictive signals of future price movements. While our research addresses the practical challenges associated with executing large orders, [Nagy et al., 2023] focuses on integrating and optimizing predictive trading signals using Q-learning methods. In contrast, [Karpe et al., 2020] highlights the dynamic and interactive nature of agent behavior and collective strategy optimization in trading environments.
V Reinforcement learning
A discrete-time Markov Decision Process (MDP) [Puterman, 1990] is defined as a tuple $\langle\mathcal{S},\mathcal{A},\mathbb{P},\mathcal{R},\gamma,\mu\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathbb{P}(\cdot|s,a)$ is a Markovian transition model that assigns to each state-action pair $(s,a)$ the probability of reaching the next state $s^{\prime}$, $\mathcal{R}(s,a)$ is a bounded reward function, $\gamma\in[0,1)$ is the discount factor, and $\mu$ is the distribution of the initial state. The policy of an agent is characterized by $\pi(\cdot|s)$, which defines for each state $s$ an action with a probability distribution over the action space.
We consider finite horizon problems in which future rewards are exponentially discounted with $\gamma$. Let us define a trajectory as a sequence of states, actions, and rewards, up to a stopping time $\tau$:
$(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},...,s_{\tau-1},a_{\tau-1},r_{\tau}).$
The objective in RL is the maximization of the expected return, given an initial state distribution:
$J_{\pi}\coloneqq\underset{s_{0}\sim\mu}{\mathbb{E}_{\pi}}\left[\sum_{i=1}^{\tau}\gamma^{i-1}r_{i}.\right].$
Where the return is the discounted sum of the rewards.
Algorithm 1 Deep Q-Network
1:Initialize replay memory $D$ to capacity $N$
2:Initialize action-value function $Q$ with random weights
3:for episode = 1 to M do
4:     Initialize state $s_{1}$
5:     for t = 1 to T do
6:         With probability $\epsilon$ select a random action $a_{t}$
7:         Otherwise select $a_{t}=\arg\max_{a}Q(s_{t},a;\theta)$
8:         Execute action $a_{t}$
9:         Observe reward $r_{t}$ and next state $s_{t+1}$
10:         Store transition ($s_{t},a_{t},r_{t},s_{t+1})$ in $D$
11:         Sample random minibatch of transitions $(s_{j},a_{j},r_{j},s_{j+1})$ from $D$
12:         Set $y_{j}=r_{j}+\gamma\max_{a^{\prime}}Q(s_{j+1},a^{\prime};\theta)$
13:         Perform a gradient descent step on $(y_{j}-Q(s_{j},a_{j};\theta))^{2}$
14:     end for
15:end for
Numerous RL algorithms are available [Schulman et al., 2017b, Schulman et al., 2015, Mnih et al., 2015]. In this work, we focus on the Deep Q-Network (DQN) algorithm, a model-free, online, off-policy reinforcement learning approach, as it provided the best learning results. The DQN algorithm is detailed in algorithm 1.
V-A Embedding Optimal Execution as an MDP
The basic building block for applying RL algorithms is a description of the environment as a MDP:
  • State: Percentage holdings remaining, percentage time remaining, volume imbalance up to 5 levels of the limit order book, best bid price, best ask price.
  • Actions: 5 possibile actions: do nothing, consume $Q_{t}^{k}=Q_{min}\times k$, $k=1,\dots,4$.
  • Reward:
    $r_{t}=\underbrace{Q_{t}^{k}\times(P_{0}-P_{t})}_{\text{implementation shortfall}}-\underbrace{\alpha d_{t}}_{\text{penalty}},$ (5)
    where $P_{t}$ is the average execution price, $P_{0}$ the arrival price, $d_{t}$ is the depth consumed by the market order at time $t$, and $\alpha>0$. Once the execution ends, even if the available time is still not over, the reward will be 0, while if the execution has not ended by the end of the execution period, there will be an additional large penalty.
The reward comprises two components: the implementation shortfall and a penalty associated with the depth consumed by the executed orders. This configuration allows the agent to develop optimal strategies for acquiring inventory while reducing significant market impact. Note that we are considering an execution problem in which we want to buy, for this reason we consider $P_{0}-P_{t}$ in the implementation shortfall. This is because if we buy at a lower price compared to the arrival price $P_{0}$, then we lower our execution costs. If we were considering a sell problem, we would need to invert to $P_{t}-P_{0}$.
V-B Tailoring the environment
Although we have the standard objective in Equation 2, how to appropriately define the MDP to achieve that objective is not straightforward. In fact, there are several factors which make that objective challenging:
  1. 1.
    The asset’s price movements are usually independent of your action. If/when the action influences the movement of the price, then it is adverse to your objective as you are generating a market impact.
  2. 2.
    It is mandatory to finish the execution by the end of the execution period, but this requirement is hard to embed in the standard reward formulation.
  3. 3.
    If you have a long execution period available, compared to the total execution size, then you can choose whether to execute rapidly, thus increasing market impact, or execute at a slower rate but risking an adverse market movement.
To tackle the first factor, we included the penalty in the reward function, which has the objective of emphasizing that we want to avoid market impact. To address the second, we include an additional large penalty if the execution is not finished by the end of the total available period. To address the third factor, the penalty in the reward function is enabling. In fact, thanks to the penalty, the reward is negative in most cases. As a reward of 0 obtained consistently once the execution is finished is preferable to a negative one reward, the learning agent should be incentivized to finish earlier.
We also explored different state and action space setups, and those described in Section V-A gave the best results.
VI Experimental Results
We defined the environment to be as realistic as possible, by using a 1 second control frequency, so that the agent has the maximum flexibility in deciding when to execute. Furthermore, we gave a long maximum execution period so as to give it the possibility to choose how much to make the execution last. In the results that follow, we focus on a buy execution, but the results hold also for a sell execution where the only thing to be changed is the sign of the implementation shortfall in the reward.
Environment setup
The environment includes several parameters and variables. The total size of the order to execute is fixed at $20000$ shares. The direction specifies whether the parent order is a buy or sell, defaulted to buy. A timeWindow of $30$ minutes is allotted for the agent to execute the entire order, with a time step duration set to $1$ second. We set the incremental size $Q_{min}=20$ for the buy or sell orders placed by the agent. Finally, the penalty is a constant amount imposed per non-executed share at the end of the time window, set at a default of $5$ per share. Additionally, a penalty of $5$ per share is also applied for over-execution beyond the intended order size. The $\alpha$ for the depth penalty in the reward is set to $2$.
RL algorithm setup
We ran a hyperparameter tuning which gave the following optimal parameters. We employed DQN, with a neural network architecture that features fully connected $2$ input layers composed of $50$ and $20$ neurons. The learning rate schedule implements a linear decrease from $10^{-3}$ to 00 in $90,000$ steps, while the exploration rate follows a $\epsilon$Greedy search from $1$ to $0.02$ in $10,000$ steps. In addition, we implement a state history length of $4$ and a market data buffer of length $50$. This approach is aimed at ensuring the stability of DQN. Finally, the objective function is a discounted sum of rewards with a discount factor of $\gamma=0.9999$.
Baseline algorithms
We utilize the following baseline algorithms to benchmark the performance of the RL policy:
  • TWAP algorithm: defined in equation 4.
  • Passive algorithm: this policy mostly keeps the agent inactive, with a 60% chance of doing nothing. Occasionally, with a 40% chance, it randomly chooses and executes a quantity among four available options, each with equal likelihood.
  • Aggressive algorithm: purchase a fixed amount $2\times Q_{\text{min}}$ each second.
  • Random algorithm: sample whether to use a market order or do nothing (quantity of 0). There is a 50% chance of doing nothing. Alternatively, it randomly selects and executes one of three actions with equal probability: do nothing, or consume $Q_{t}^{k}=Q_{min}\times k$, with $k=1,\dots,3$.
VI-A Experiments
Next, we evaluate the RL agent’s ability to execute large orders while minimizing market impact. This involves a detailed comparison of the RL agent’s performance with the baseline strategies mentioned earlier.
Figure 3: Average episode reward for DQN agent with a environment seed equal to 10 and learning rates equal to $10^{-2}$ (pink), $10^{-3}$ (red), and $10^{-4}$ (blue).
Figure 3 shows the learning curve of the DQN algorithm. The increasing reward curve indicates that the environment is correctly formulated and that the agent is maximizing implementation shortfall and reducing market impact during training, approaching a zero average reward as it learns.
Figure 4: Implementation shortfall normalized by the initial order size distribution.
In Figure 4, we evaluate out-of-sample the optimized reinforcement learning policy and compare it to the three baselines. The distribution is created after running 20, $30$-minute windows with different seeds. We observe the average step implementation shortfall (see Equation 5). In this case we see that the RL agent generates a lower variance compared to the baselines with also a higher average. This can be interpreted as the fact that the RL execution strategy is capable of executing consistently close to the arrival price while minimizing market impact.
Figure 5: Ask prices during the RL execution.
Figure 5 presents the ask prices during execution by the RL agent over time. The stability of ask prices under the RL policy indicates the agent’s ability to minimize market impact and adverse price movements effectively.
Figure 6: Execution trajectory of the RL agent: amount executed over time.
Figure 6 shows the execution trajectory of the RL agent: how much it has executed of the total size. The agent executes a significant portion of its holdings rapidly, followed by a more controlled and steady approach as the execution period progresses. This pattern aligns with optimal execution principles, balancing the trade-off between immediate market impact and long-term price stability. By front-loading the execution and then tapering off, the agent minimizes the risk of adverse market movements while efficiently managing its inventory.
Figure 7: Bid and ask prices along with actions taken by the RL agent over time.
Figure 7 provides a detailed view of the bid and ask prices alongside the actions executed by the RL agent at various time steps. The figure illustrates that the actions are aligned with a minimal deviation between the bid and ask prices, indicating effective control over the execution strategy without causing market disruption.
Figure 8: Distribution of the spreads generated by the different strategies.
In Figure 8, we observe the distribution of price spreads generated by the different trading strategies. We can see that the RL widens a bit the spreads compared to the TWAP, but this is because it finishes the execution earlier as we saw in Figure 6. If we consider instead the aggressive strategy, it generates very wide spreads and so was taken out of the plot in order to visualize better the other policies.
Figure 9: Distribution of the volume imbalance generated by the different strategies.
Figure 9 shows the distribution of the volume imbalance generated by the RL agent’s execution. A market impact would be visible with a skewed imabalance. In this case the imbalance is centered, indicating a consistent strategy for maintaining balance in the LOB.
Metric RL TWAP Passive Random
Imp. Short. -0.022 -0.028 -0.024 -0.049
Avg. Pen. -0.001 $0$0.0035 -0.003
Time % 0.451 0.998 0.727 0.553
TABLE I: Comparison of normalized performance metrics across different trading strategies.
Table I summarizes some of the metrics also present in the plots, adding also the aggressive strategy. The first two rows represent the two parts of the reward defined in Equation 5. The last row shows on average how much time is necessary to finish the execution.
The RL agent’s ability to keep the order book balanced can be attributed to its adaptive response to real-time market conditions. When encountering a growing imbalance, the agent likely adjusts its trading strategy to prevent further imbalance, thus avoiding significant deviations in asset prices. By maintaining smaller imbalances in the order book, the RL agent reduces the risk of market impact, which often leads to unfavorable price movements and increased transaction costs.
VII Conclusions
The results of this study demonstrate the effectiveness of using reinforcement learning (RL) for optimal execution in financial markets. Using the deep Q-Network (DQN) algorithm within a simulated market environment, we observed significant improvements in execution performance compared to the chosen benchmark strategies.
Our experiments show that the RL agent consistently achieved higher returns and lower variance in implementation shortfall. The agent’s ability to adapt to market conditions and execute trades close to the arrival price resulted in minimized market impact and transaction costs. This stability is crucial to maintain favorable execution conditions.
The execution trajectory analysis highlighted the strategic balance of the RL agent between immediate market impact and long-term price stability. The agent effectively managed its holdings, executing a significant portion rapidly at the beginning and then proceeding more steadily. This approach is aligned with the principles of optimal execution, ensuring efficient inventory management.
Although our findings are promising, it is important to acknowledge areas for future exploration to enhance the robustness and applicability of RL-based strategies. The simulated market environment, though realistic, can be further refined to capture a wider array of market dynamics and participant behaviors. Expanding the range of market conditions under which the RL models are tested will provide deeper insights into their performance and adaptability.
Furthermore, the computational resources required for training RL models are considerable, and optimizing these processes will be crucial for practical implementation. Streamlining the training process and improving the efficiency of the algorithms will facilitate their broader application in real-world scenarios.
In conclusion, the positive results of our study underscore the potential of reinforcement learning to address the complexities of financial markets. This approach provides traders and institutions with enhanced tools for efficient and effective execution. Continued research and development will be essential to ensure the robustness and adaptability of these models, paving the way for their successful application in dynamic and competitive trading environments.